Write the equation for the circle with center at (0,0) and radius of 4.
Answer by nerdybill(7384)

Standard form of a circle:
(x−h)^2+(y−k)^2=r^2
.
plug in center at (0,0) and radius of 4
(x−0)^2+(y−0)^2=4^2
x^2+y^2 = 16 (this is what they're looking for)